The problem I have is
7-3(x+3)=-5(x-2)+2x-12

Answer by fanism(39)

this has no solution, the unknown cancelled out.
check your question!
7-3(x+3)=-5(x-2)+2x-12
7 - 3x - 9 = -5x + 10 + 2x - 12
-2 - 3x = -3x - 2
0 = 0